OpenStudy (anonymous):

Tina bought a coat that cost $70.00. The sales tax in her state is 6%. What is the final amount that Tina has to pay for the coat? i think it is $74.20

OpenStudy (anonymous):

@katherinkoon

OpenStudy (anonymous):

@rebeccaxhawaii am i right? this is my last one

rebeccaxhawaii (rebeccaxhawaii):

yes

rebeccaxhawaii (rebeccaxhawaii):

that is correct

OpenStudy (anonymous):

i got a 62 :|

rebeccaxhawaii (rebeccaxhawaii):

how ?

OpenStudy (anonymous):

idk thanks for trying tho

OpenStudy (popeyes):

From what I got, you are correct. Turn 6% into a decimal: 6/100 = 0.06 Multiply: 0.06 * 70.00 = 4.20 Add: 70.00 + 4.20 = 74.20 So, the answer is $74.20, like you said.